Controlling Interest in Income
On January 1, 2014, Sherwood Company, an 80% owned subsidiary of Paradise Company, sold to Paradise
Company equipment with a book value of $600,000 for $840,000. The equipment had an estimated remaining
useful life of eight years on the date of the intercompany sale.
Paradise Company reported net income from its independent operations of $550,000, and Sherwood
Company reported net income of $300,000 in the years of 2014 and 2015.
Required:
Calculate the controlling interest in combined net income for the years ended December 31, 2014, and
December 31, 2015.
Answer #1
Particulars $2,014 $2,015
Independent income of paradise $550,000 $550,000
Less : Depreciation $30,000 $30,000
Independent income of paradise $520,000 $520,000
Income of Sherwood $300,000 $300,000
Less Profit on Sale of Equipments $240,000                        -  
Income $60,000 $300,000
Less Minority Shareholding@80% $48,000 $240,000
Combined Income $568,000 $760,000

Minority Shareholding is the percentage of shareholding hold by entity other than holding company.

Workings
Book Value of Equipment Sold $600,000
Sold Value of Equipments $840,000
Profit on Sale of Equipments $240,000
Useful Life of Assets 8 years
Depreciation on assets
On Book Value $75,000 =600000/8
On Sold Value of Equipments $105,000 =840000/8
Difference is Depreciation $30,000 =105000-75000